Grade 11Physical Chemistry

(i) if mass of an electron is 9.11×10^-31 kg how many electrons would weigh 1 kg?(ii) how many astronomical units are there in 1 metre?

Shalin
9 Years agoGrade 11

1 Answer

Neeraj patel
9 Years ago
Best on unitary method u can deal both part of your question comfortably​ .for (i) 9.11×10^-31kg=1e 1kg=(1/9.11×10^-31)e =10^30electrons (approx)For (ii) 1.496×10^11m=1Au 1m=(1/1.496×10^11)Au =6.685×10^-12Au